OpenStudy (jayde742):

Let f(x) = -3x - 4 and g(x) = 2x - 6. find ( F*g) (-7) A. 28 B. -20 C. 17 D. 56

OpenStudy (jayde742):

@agent0smith

OpenStudy (agent0smith):

Plug -7 into f(x) Plug -7 into g(x) Multiply those two results

OpenStudy (jayde742):

but i dot get it ugh every time i do it it comes out completely wrong

OpenStudy (agent0smith):

Show your work

OpenStudy (jayde742):

well i think im setting the whole problem up wrong i dont know how to set it up

OpenStudy (jayde742):

f (-7) = -3x -4 amd g (-7) = 2x -6?

OpenStudy (agent0smith):

You didn't plug in the -7 for x...

satellite73 (satellite73):

\[ f(x) = -3x - 4\\ f(-7) = -3\times (-7) - 4\]

OpenStudy (triciaal):

-7 in g(x) then that value in f(x) so 56